# Temperature-dependent Raman spectra of 2H-MoS2 from Machine Learning-driven statistical sampling

### 02 Apr 2026 v1

@@ III Results and discussion
In this section we first evaluate the performance of the MLIP, then we present the the results of TDEP-DFPT approach for the calculation of the Raman spectrum with both MD-based and stochastic sampling. Our analysis focuses on the evolution in temperature of the Raman active modes A1g and E2g, with particular attention to phonon frequency shift and spectral broadening.
@@ III.1 The MLIP training
We train the MTP model with the following hyperparameters (see Ref. [50]): {level: 20, cutoff radius: 8 Å, radial basis size: 8, radial basis type: Chebyshev polynomials; weights: (1,1,1)}. For the k-fold cross-validation the average root mean squared errors (rmse) on energy, forces and stress are below 0.003 eV/atom, 0.022 eV/Å and 0.0005 eV/Å2, respectively, with a dataset size of 339 structures with 96 atoms each. The final training (with the same hyperparameters) is performed on the entire dataset and achieves rmse and R2 of 0.00176 eV/atom, 0.01796 eV/Å and 0.00035 eV/Å2 on the energy, forces and stress, respectively. As a further confirm of our MLIP being able to align to DFT, we use the MLIP to perform a structure relaxation with LAMMPS. The optimized cell parameters differ by less than 0.00012 Å and 0.027 Å for $a$ (and $b$) and $c$, respectively.
@@ III.2 Thermalized unit cells
We determine the lattice parameters of the thermalized unit cell between 100 and 700 K, according to the method described in Sect. II.2.1, using LAMMPS [62] for the MD simulations, as in the rest of the study. In Fig. 1 the lattice parameter variation with temperature is shown, comparing to experimental measurements from Ref. [13, 45]. Our calculated parameters are in excellent agreement with the experimental ones, overestimating by only $\sim$$0.16\%$ and $\sim$$0.2\%$ respectively for $a$ and $c$, which is within the typical accuracy of GGA calculations with vdW corrections for this kind of layered systems [22, 14]. A good estimation of the lattice constants is crucial to ensure that the model for the PES, and therefore any derived quantity, is meaningful.
@@ III.3 Vibrational properties
We use TDEP to extract the IFC2s and IFC3s. The phonon dispersion bands vs. T are reported in Fig. 2. As expected, thermal expansion and anharmonicity leads to a general softening of the modes.
In Fig. 3 we show our phonon bands obtained at 300 K with both the MD and stochastic sampling, as compared to the experimental values from Ref. [64]. We note that the two sampling methods give very similar phonon bands. Our frequencies are smaller than the experimental ones, by a factor which increases with the frequency itself. We attribute this to the DFT approximations used to train the MLIP, as GGA functionals systematically underestimate the phonon frequencies. This discrepancy is also consistent with the slight overestimation of the lattice constants we present in the previous section.
In agreement with experiments [69, 74, 58, 63, 33, 36, 72, 70] and symmetry, the two degenerate E2g in-plane (x-y) modes and the A1g out-of-plane (z) mode are found to be Raman active in 2H-MoS2. Two other E2g degenerate modes are present at low frequency, but they are associated to sheer displacement of the layers. Due to their low effect on the electronic polarizability, they show a Raman response orders of magnitude smaller than the first two, and often noisy. Therefore, in this study we focus on the high-frequency modes.
@@
We start by noting that for both modes a temperature increase leads to a red shift in the frequencies. This is the expected behavior as the unit cell expands as a result of increasing temperature. We fit the temperature-evolution with a linear function (intercepts, coefficients and experimental references are reported in Table 1). Looking at the trends of the frequencies in the full range of temperature, we note that the MD sampling is more sensitive to temperature. Since the same unit cell is used for both the MD and stochastic sampling (the first-order optimizations performed in the second method leads to negligible changes), the difference in the rate of change of the frequencies can simply be ascribed to the quantum phonon occupation in the stochastic method, which starts larger and increases less rapidly in temperature than in the classical case. The rate of change in the occupation is related to that of the mean square displacement which determines how much more anharmonicity is captured upon a temperature increase.
@@
Our widths tend to be smaller than those observed experimentally. This can be ascribed to higher-order interactions beyond the third order, and to disorder, inhomogeneous broadening, and measurement noise in the experimental FWHMs. The latter factors complicate a strict quantitative comparison with the purely anharmonic linewidths calculated here. However, our widths follow a similar trend vs. T as in the experiments. We also note that the stochastic sampling systematically leads to larger FWHMs than the MD sampling.
@@
A further analysis of the error, e.g. by including the fourth-order terms, would be beneficial to understand the evolution of the anharmonic properties versus temperature. While with TDEP it is possible to calculate the imaginary part of the self energy at the fourth order evaluated at the harmonic phonon frequency, $\Gamma^{(4)}(\omega_{\lambda})$, the calculation of the full fourth-order frequency-dependent self energy is very involved and not implemented yet. Furthermore, the real counterpart $\Delta^{(4)}(\omega_{\lambda})$ is not calculated either. Therefore, as a first inspection, we calculated the fourth-order contribution to the self-energy, to include it as a rigid translation of the imaginary part:
@@
with $A$, $2\Gamma$, and $\tilde{\omega}$ being the area, the FWHM, and the center of the peak, respectively.
In Fig. 7 we plot the ratio between the averaged Raman intensity of the E2g and A1g modes. The ratio between our Raman intensities are of the same order of magnitude as the experimental ones, which, unlike ours, exhibit an overall decreasing non-monotonic trend. Our calculated spectra (Fig. 6) show higher E2g peaks than the A1g. While this aligns with some experimental reports [15, 47, 48], it contradicts others [18, 35]. This discrepancy likely stems from the fact that relative intensities are highly sensitive to the exciting frequency. In fact, any frequency above the 1.23 eV band gap [29], can easily trigger resonant Raman phenomena, while our calculations are performed in the static limit. This highlights the difficulty in comparing different experiments with theory or amongst themselves. For intensities and peak ratios the conditions, substrate, dielectric environment, and laser excitation frequency and polarization must be known very precisely.
@@ IV Conclusions
We calculate the Raman spectra of 2H-MoS2 from 100 to 700 K with the TDEP approach, so as to include finite-T effects on the position and width of the Raman peaks. We generate a representative dataset and use it to train an accurate MTP model, which accelerates the determination of the thermalized cell parameters, and the thermodynamic sampling. We compare two different sampling methods: classical MD simulation of the atomic trajectories through phase space, and stochastic generation of structures from an effective canonical ensemble according to thermally renormalized phonons. We obtain cell parameters in very good agreement with experimental measurements. Our phonon band dispersions at room temperature also agree with those from experiments, confirming our method gives a good description of the lattice dynamics at the harmonic level, including the renormalization coming from the interaction between phonons. We calculate the phonon self-energy with third-order contributions, obtaining phonon lifetimes which compare well to the experimentally observed ones, in absolute terms and for their temperature dependence. These results could be improved by accounting for contributions of phonon processes at higher-order in a complete way (full frequency-dependent phonon self-energy at the fourth order), as well as electron-phonon interactions and electronic screening modulations due to many-body effects. The Raman intensities obtained from the DFPT dielectric calculations are compared with those extracted from the experimental spectra. They are of similar order of magnitude, but depend very strongly on the experimental (and numerical) conditions. We ascribe at least part of this discrepancy to most measurements being performed in the resonant Raman regime. Future experimental works with exciting light below the band gap might confirm (or contradict) our resonance-free results.

### 09 Apr 2026 v2

@@ III Results and discussion
In this section we first evaluate the performance of the MLIP, then we present the results of the TDEP-DFPT approach for the calculation of the Raman spectrum with both MD-based and stochastic sampling. Our analysis focuses on the evolution in temperature of the Raman active modes A1g and E2g, with particular attention to phonon frequency shift and spectral broadening.
@@ III.1 The MLIP training
We train the MTP model with the following hyperparameters (see Ref. [50]): {level: 20, cutoff radius: 8 Å, radial basis size: 8, radial basis type: Chebyshev polynomials; weights: (1,1,1)}. For the k-fold cross-validation, the average root mean squared errors (rmse) on energy, forces and stress are below 0.003 eV/atom, 0.022 eV/Å and 0.0005 eV/Å3, respectively, with a dataset size of 339 structures with 96 atoms each. The final training (with the same hyperparameters) is performed on the entire dataset and achieves rmse and R2 of 0.00176 eV/atom, 0.01796 eV/Å, and 0.00035 eV/Å3 on the energy, forces and stress, respectively. As a further confirmation of our MLIP being able to align to DFT, we use the MLIP to perform a structure relaxation with LAMMPS [62]. The optimized cell parameters differ by less than 0.00012 Å and 0.027 Å for $a$ (and $b$) and $c$, respectively.
@@ III.2 Thermalized unit cells
We determine the lattice parameters of the thermalized unit cell between 100 and 700 K, according to the method described in Sect. II.2.1, using LAMMPS for the MD simulations, as in the rest of the study. In Fig. 1 the variation in the lattice parameters with temperature is shown, compared to the experimental measurements from Ref. [13, 45]. Our calculated parameters are in excellent agreement with the experimental ones, overestimating by only $\sim$$0.16\%$ and $\sim$$0.2\%$ respectively for $a$ and $c$, which is within the typical accuracy of GGA calculations with van der Waals corrections for this kind of layered systems [22, 14]. A good estimation of the lattice constants is crucial to ensure that the model for the PES, and therefore any derived quantity, is meaningful.
@@ III.3 Vibrational properties
We use TDEP to extract the IFC2s and IFC3s. The phonon dispersion bands vs. T are reported in Fig. 2. As expected, thermal expansion and anharmonicity lead to a general softening of the modes.
In Fig. 3 we show our phonon bands obtained at 300 K with the MD and stochastic sampling, as compared to the experimental values from Ref. [64]. We note that the two sampling methods give very similar phonon bands. Our frequencies are smaller than the experimental ones by a factor which increases with the frequency itself. We attribute this to the DFT approximations used to train the MLIP, as GGA functionals systematically underestimate the phonon frequencies. This discrepancy is also consistent with the slight overestimation of the lattice constants that we present in the previous section.
In agreement with experiments [69, 74, 58, 63, 33, 36, 72, 70] and symmetry, the two degenerate E2g in-plane (x-y) modes and the A1g out-of-plane (z) mode are found to be Raman active in 2H-MoS2. Two other E2g degenerate modes are present at low frequency, but they are associated to shear displacement of the layers. Due to their low effect on the electronic polarizability, they show a Raman response orders of magnitude smaller than the first two, and often noisy. Therefore, in this study we focus on the high-frequency modes.
@@
We start by noting that for both modes a temperature increase leads to a red shift in the frequencies. This is the expected behavior as the unit cell expands as a result of increasing temperature. We fit the temperature-evolution with a linear function (intercepts, coefficients and experimental references are reported in Table 1). Looking at the trends of the frequencies in the full range of temperature, we note that the MD sampling is more sensitive to temperature. Since the same unit cell is used for both the MD and stochastic sampling (the first-order optimizations performed in the second method lead to negligible changes), the difference in the rate of change of the frequencies can simply be ascribed to the quantum phonon occupation in the stochastic method, which starts larger and increases less rapidly in temperature than in the classical case. The rate of change in the occupation is related to that of the mean square displacement which determines how much more anharmonicity is captured upon a temperature increase.
@@
Our widths tend to be smaller than those observed experimentally. This can be ascribed to interactions beyond the third order, and to disorder, inhomogeneous broadening, and measurement noise in the experimental FWHMs. The latter factors complicate a strict quantitative comparison with the purely anharmonic linewidths calculated here. However, our widths follow a similar trend vs. T as in the experiments. We also note that the stochastic sampling systematically leads to larger FWHMs than the MD sampling.
@@
A further analysis of the error, e.g. by including the fourth-order terms, would be beneficial to understand the evolution of the anharmonic properties versus temperature. While with TDEP it is possible to calculate the imaginary part of the self energy at the fourth order evaluated at the harmonic phonon frequency, $\Gamma^{(4)}(\omega_{\lambda})$, the calculation of the full fourth-order frequency-dependent self-energy is very involved and not implemented yet. Furthermore, the real counterpart $\Delta^{(4)}(\omega_{\lambda})$ is not calculated either. Therefore, as a first inspection, we calculated the fourth-order contribution to the self-energy, to include it as a rigid translation of the imaginary part:
@@
with AA, $2\Gamma$, and $\tilde{\omega}$ being the area, the FWHM, and the center of the peak, respectively.
In Fig. 7 we plot the ratio between the averaged Raman intensity of the E2g and A1g modes. The ratios between our Raman intensities are of the same order of magnitude as the experimental ones, which, unlike ours, exhibit an overall decreasing non-monotonic trend. Our calculated spectra (Fig. 6) show higher E2g peaks than the A1g. While this aligns with some experimental reports [15, 47, 48], it contradicts others [18, 35]. This discrepancy likely stems from the fact that relative intensities are highly sensitive to the exciting frequency. In fact, any frequency above the 1.23 eV band gap [29], can easily trigger resonant Raman phenomena, while our calculations are performed in the static limit. This highlights the difficulty in comparing different experiments with theory or amongst themselves. For intensities and peak ratios the conditions, substrate, dielectric environment, and laser excitation frequency and polarization must be known very precisely.
@@ IV Conclusions
We calculate the Raman spectra of 2H-MoS2 from 100 to 700 K with the TDEP approach, so as to include finite-T effects on the position and width of the Raman peaks. We generate a representative dataset and use it to train an accurate MTP model, which accelerates the determination of the thermalized cell parameters, and the thermodynamic sampling. We compare two different sampling methods: classical MD simulation of the atomic trajectories through phase space, and stochastic generation of structures from an effective canonical ensemble according to thermally renormalized phonons. We obtain cell parameters in very good agreement with experimental measurements. Our phonon band dispersions at room temperature also agree with those from experiments, confirming that our method gives a good description of the lattice dynamics at the harmonic level, including the renormalization coming from the interaction between phonons. We calculate the phonon self-energy with third-order contributions, obtaining phonon lifetimes which compare well to the experimentally observed ones, in absolute terms and for their temperature dependence. These results could be improved by accounting for contributions of phonon processes at higher-order in a complete way (full frequency-dependent phonon self-energy at the fourth order), as well as electron-phonon interactions and electronic screening modulations due to many-body effects. The Raman intensities obtained from the DFPT dielectric calculations are compared with those extracted from the experimental spectra. They are of similar order of magnitude, but depend very strongly on the experimental (and numerical) conditions. We ascribe at least part of this discrepancy to most measurements being performed in the resonant Raman regime. Future experimental works with exciting light below the band gap might confirm (or contradict) our resonance-free results.